Two coins are tossed. Let T be the number of tails that occurs. Determine the values 

of the random variable T.


1
Expert's answer
2021-02-24T07:50:51-0500

When two coin are tossed, the sample space is {"HH,HT,TH,TT" }

Total number of samples ="4"


Let "\text{X}" represent the number of tails.

Therefore, "\text{X}" can take the value of "0,1 \text{ or } 2." .


So, Probability of getting 0 tails

"P(X=0)=\dfrac{1}{4}"


Probability of getting 1 tails

"P(X=1)=\dfrac{2}{4}=\dfrac{1}{2}"


Probability of getting 2 tails

"P(X=2)=\dfrac{1}{4}"


 Therefore The Probability distribution is given by-